OpenStudy (anonymous):

find the limit of lim x tends to 0^-[e^(1/x)]

OpenStudy (anonymous):

not clear. limit as x goes to what? limit of what?

OpenStudy (anonymous):

\[lim_{x\rightarrow 0}e^{\frac{1}{x}}\]?

OpenStudy (anonymous):

correct but x tends to 0^-1 (power of negative 1)

OpenStudy (anonymous):

Your teacher is trying to be cute, or tricky. 0^-1=1/0=infinity. So find the limit as x goes to infinity.